# Evolving Large Language Model Assistant with Long-Term Conditional Memory

### v1

@@ Introduction
Table 6: The GPT-score and GSB result for the ablation study. The GSB test compare the ablated models and the original ones.
We summarize the experiment result with the following conclusions.
Conditional memory achieves the relatively best result among the three forms of memory. In Table 4 and Figure 5, we display the effectiveness of different forms of memory on all three datasets. Our proposed conditional memory achieves the best result in learning new knowledge and learning from human feedback, which shows the ability of conditional memory to abstract and retain important information from the dialogue. As for continuing previous dialogue, the best result is achieved by directly using dialogue history as memory, and the other two forms of memory also obtain a comparable result. Considering that this task usually requires more details, using dialogue history can have more advantages.
Multi-view memory does not always brings better performance, but the combination between conditional memory and summary-based memory can improve the performance. In Table 3, we show the influence of combining conditional memory with the other two forms of memory as multi-view memory. On the one hand, we can observe that history-based memory leads to a performance drop in learning new knowledge and learning from human feedback and a small improvement in continuing previous dialogue. This suggests that there is a overlap between history-based memory and conditional memory and the combination usage of them brings redundancy rather than useful information. On the other hand, we find that summary-based memory is an effective complement to conditional memory and brings improvement in all datasets.
Self-reflection retrieval is effective in most cases, especially on summary-based memory. To evaluate the effectiveness of self-reflection retrieval, we apply it to the LLM assistant with all three types of memory. Here, we limit the self-reflection to no more than one time. As shown in Table 5, in most cases, adding a self-reflection retrieval makes the retrieval results more accurate and brings a slight improvement in the outcomes. It is particularly evident in the one with summary-based memory. We believe this is because summary-based memory contains far more information than is required for user input, resulting in relatively poor search accuracy. And the self-reflection retrieval can alleviate this situation.
@@
In Figure 6, we conduct a brief analysis of the effect of the number of retrieved memory records. For the fine-grained memory records, history-based memory and our conditional memory, the result is no longer significantly improved after $K$ reached 3. Hence, we select the top-3 memory records for history-based memory and conditional memory in our experiment. For the coarse-grained summary-based memory records, considering information related to the user input only exists within a single dialogue in most cases, the value of $K$ has little effect on the result. Therefore, we only select one memory record for summary-based memory.
Conclusion
@@
In this paper, we propose to use long-term conditional memory to build an easy-access evolving Large Language Model assistant that can learn experience or knowledge during its dialogue with the human user. To investigate the model in our scenario, we focus on the influence of different forms of memory records and the effective usage of the retrieved memory records in generating the response. A novel approach is proposed to construct conditional memory. Moreover, considering the lack of dataset for evaluation of the framework, we also build three test datasets based on different abilities required by an evolving LLM assistant. The experiments further show the effectiveness of our proposed methods and bring some interesting conclusions. It is worth pointing out that we only investigate the foundation of the idea in this paper, and other key points such as the time stamp or forgetting mechanism are still waiting to be explored in further works.
Appendix A Details for Method
Prompt for Summary-Based Memory Construction
@@ Appendix A Details for Method
Dialogue summaries can also be used as memory. We use LLM to generate a summary for each dialogue session. The instruction prompt for generating summaries for the dialogue is shown in Figure 7.
Figure 7: Prompt for summary-based memory construction.
Prompt for Conditional Memory Construction
@@
The construction of conditional memory mainly consists of two parts: the decision-making together with the response generation and the memory construction. We display the prompt for the first part in Figure 8. Here, user input is the input question or instruction from the user and memory stands for the retrieved memory records. It is worth noticing that the context for the current dialogue is not in the prompt, but it is entered as a dictionary like normal LLM assistant. Compared with the normal response generation, we can see that we convert it to a two-step instruction and give an output example for it to better separate the outcome. Only the generated response will be shown to the user and the yes/no decision is hidden behind the model.
@@
Figure 8: Prompt for decision making in conditional memory construction.
@@
For the second part of memory construction, we display its prompt in Figure 9. Here, the dialogue context refers to ongoing dialogue just including the current utterance. It is also decomposed into three steps. The first step requires the LLM to find the related context from all dialogue contexts. The second step requires the LLM to summarize the related context. And the last step requires the LLM to generate knowledge that is worth memorizing. For the output, we only store the summary and knowledge as the memory record, and the related context is considered to be a mid product.
Figure 9: Prompt for memory record construction in conditional memory construction.
Prompt for Self-Reflection Retrieval
@@
This self-reflection retrieval contains two steps and we show its prompt in Figure 10. The input includes the user input and a set of retrieved candidate memory records. We ask the LLM to make a decision about whether the retrieved information is enough for it to respond to the user input. If it is enough, the LLM assistant selects the related information from these memory records as output. If it is not enough, the LLM assistant is required to improve the original query (user input) to find more information. In practice, the improvement is reified as generating some keywords/phrases representing the missing information.
Figure 10: Prompt for self-reflection retrieval.
Appendix B Details for Datasets Construction
Prompt for Continuing Previous Dialogue
When constructing the test dataset for continuing the previous dialogue, we aim to continue writing a existing dialogue. The related prompt is shown in Figure 11.
Figure 11: Prompt for constructing continuing previous dialogue dataset.
Prompt for Learning New Knowledge
For constructing the test dataset for learning new knowledge, we need to transfer a set of personal settings to a question-answering test utterance. Considering it is a relatively easy task, we adopt in-context learning to solve this problem. We display the in-context examples we used in Figure 12.
Figure 12: Prompt for constructing learning new knowledge dataset.
Prompt for Learning from Human Feedback
To construct the test dataset for learning from human feedback, we decompose it into three steps with different instruction prompts. Given a seed question/instruction, the first step is to generate a similar question/instruction used for testing. Here, we also use in-context learning. The in-context examples are shown in Figure 13. Then we use the LLM to obtain an initial response for the two questions/instructions. The second step is to generate feedback that can be applied to the two both input-response pairs. Hence, we take both pairs as input and use the LLM to generate feedback for them. The instruction prompt is displayed in Figure 14. With the feedback, we can get an optimized response from the LLM for both seed input and test input. To complete the source dialogue, the last step is to generate a background dialogue for this input-respond-feedback-respond process. The related instruction prompt is shown in Figure 15. Here, the input in the prompt is the seed question/instruction.
Figure 13: Prompt for step 1 of constructing learning from human feedback dataset.
Figure 14: Prompt for step 2 of constructing learning from human feedback dataset.
Figure 15: Prompt for step 3 of constructing learning from human feedback dataset.
Appendix C Details for GPT Evaluation
Prompt for Scoring (GPT-score)
GPT-score aims to give a predicted response a score between a predefined margin. The LLM is required to give a 0-2 score for a response. Here, “0” represents the response is totally wrong, “1” represents the response contains part of personal information in the reference and “2” represents the response contains all information. We display the related prompt in Figure 16. The input of the prompt is the reference response and the predicted response.
Figure 16: Prompt for scoring (GPT-score).
Prompt for Comparing (GSB test)
Comparison aims to conduct a Good/Same/Bad test (GSB) between two candidates. The LLM is used to make a decision about which candidate is better given the reference response and other necessary information. As shown in Figure 17, the two candidates are assigned to option A or option B, and we also provide the user input ([Question] in the prompt) and the reference response for the background information.
Figure 17: Prompt for comparing (GSB test).
Prompt for Multiple Choice
Multiple choice is similar to comparison, but the LLM needs to choose the best candidate among A, B, C, and D, four options. The corresponding prompt is shown in Figure 18.
Figure 18: Prompt for Multiple Choice.

### v2

@@ 5.3 Experiment Results
We summarize the experiment result as the following conclusions.
Conditional memory achieves the relatively best result among the three forms of memory. In Table 4 and Figure 5, we display the effectiveness of different forms of memory on test datasets. The conditional memory achieves the best result in learning new knowledge and learning from human feedback, which shows the ability of conditional memory to abstract and retain important dialogue information. As for continuing previous dialogue, the best result is achieved by directly using dialogue history as memory, and the other two forms of memory also obtain a comparable result. Considering that this task usually requires more details, using dialogue history can have more advantages.
Multi-view memory does not always brings better performance, but the combination between conditional memory and summary-based memory can improve the performance. In Table 3, we show the influence of combining conditional memory with the other two memory types as the multi-view memory. On the one hand, we find that history-based memory leads to a performance drop in learning new knowledge and learning from human feedback and a small improvement in continuing previous dialogue. This suggests that there is a overlap between history-based memory and conditional memory and the combination usage of them brings redundancy rather than useful information. On the other hand, we find that summary-based memory is an effective complement to conditional memory and brings improvement in all datasets.
@@
Table 6: The GPT-score and GSB result for the ablation study. It compares the ablated models and the original ones.
Self-reflection retrieval is effective in most cases, especially on summary-based memory. To evaluate the effectiveness of self-reflection retrieval, we apply it to the LLM assistant with all three types of memory. Here, we limit the self-reflection to no more than one time. As shown in Table 5, in most cases, adding a self-reflection retrieval makes the retrieval results more accurate and brings a slight improvement in the outcomes. It is particularly evident in the one with summary-based memory. Considering summary-based memory contains far more information than is required for user input, it leads to relatively poor search accuracy. And the self-reflection retrieval can alleviate this situation.
@@ 5.4 Analysis and Discussion
In Figure 6, we conduct a brief analysis of the effect of the number of retrieved memory records. For the fine-grained memory records, history-based memory and our conditional memory, the result is no longer significantly improved after $K$ reached 3. Hence, we select the top-3 memory records for history-based memory and conditional memory in our experiment. For the coarse-grained summary-based memory records, the value of $K$ has little effect on the result. Hence, we select one memory record for summary-based memory. Combining with memory records length in Table 1, adding conditional memory to the assistant increase 140 tokens to the prompt size per utterance on average. This is far more smaller than dialogue history with thousands of tokens, which helps alleviate the problems from having a limited context window.
6 Conclusion
@@
In this paper, we propose to use evolving conditional memory to build an easy-access personalized Large Language Model assistant that can learn experience or knowledge during its dialogue with the human user. We focus on the influence of different forms of memory records and the effective usage of the retrieved memory records in generating the response. Moreover, considering the lack of dataset for evaluation in such scenario, we build three test datasets based on different abilities required by a personalized LLM assistant. The experiments show the effectiveness of our proposed methods and bring interesting conclusions.
Limitations
One limitation for this work is that we only use GPT-4 rather than other LLM (GPT-3.5, Vicuna) in this paper. And the main reason we use GPT-4 is that it is the strongest LLM now. And we believe this task and the experiments are not easy. So we would like to test the effectiveness of different memory mechanism without the limitation of the LLM. Using weaker model like GPT-3.5 may lead to more results where no model can provide the right answer. But we do think that testing on different LLMs can provide more insights. Meanwhile, it is worth pointing out that we only investigate the foundation of the idea in this paper, and other key points such as the time stamp or forgetting mechanism are still waiting to be explored in further works.
Ethics Statement
The scientific artifacts we used are available for research with permissive licenses, including ROUGE and the GPT-4 model. The use of these artifacts is consistent with their intended use. Considering all the datasets we constructed are built up based on GPT-4 and other existing public available datasets, we think there are no potential ethics risks. All in all, this work complies with the ACL Ethics Policy.
Appendix A Details for Datasets Construction
A.1 Human Evaluation for the Constructed Datasets
To better understand the quality of the constructed testing dataset, we conduct a small-scale sanity check by human annotators. For each dataset, we randomly select 50 data samples and employ three fluent English speakers to check each of them. The human annotators need to answer two question: (1) Whether the test utterance is related to the source dialogue; (2) Whether the respond in the testing utterance is appropriate based on the source dialogue. If the answers for both questions are yes, the data point is considered as qualified.
As shown in the Table 7, we display the qualified percentage of data samples and the Fleiss Kappa coefficient between the three human annotators. We can find that the qualified rate for Learning New Knowledge dataset is relatively lower. The reason is that the dataset is constructed based on multi-session chat (MSC) dataset. In MSC, a set of personal information is firstly given and the annotators are required to conduct the multi-session chat based on these information. In our dataset, we just directly transfer these personal information into question-knowledge pairs for testing. So it is possible that not all personal information are mentioned in the dialogue in MSC dataset. Considering there is over 1000 testing utterance in Learning New Knowledge dataset, these unqualified data sample will not largely effect the experiment result. After all, no model can provide a right answer when facing these testing utterance.
Dataset Qua Kap
Learning New Knowledge 60% 0.676
Learning from Feedback 84% 0.746
Continuing Previous Dialogue 81% 0.492
Table 7: Human Evaluation for the Constructed Datasets. Here, “Qua" stands for the qualified percentage of data samples, and “Kap" refers to the Fleiss Kappa coefficient between the three human annotators.
A.2 Prompt for Continuing Previous Dialogue
When constructing the test dataset for continuing the previous dialogue, we aim to continue writing a existing dialogue. The related prompt is shown in Table 8.
————–Input————- [Dialogue]: User: ….. Assistant: ….. ….. User: ….. Assistant: ….. ————–Prompt————- The input is a [Dialogue] between a human user and a AI assistant. The background situation is that after the [Dialogue] is finished, some time later, the user want to continue this dialogue with the assistant. Following the previous dialogue, you are reuqired to extend the dialogue to one more turn (user starts the first and then assistant make the respond). The new dialogue turn should follow closely with the [Dialogue]. The new dialogue turn must to utilize or mention the information from the [Dialogue]. In other words, the new dialogue turn can not be independent from the [Dialogue]. The user input can not be respond without the [Dialogue] or the respond is different with or without the [Dialogue]. But do not use pronoun to represent any information in [Dialogue] for the new dialogue. As for the user in the new dialogue, do not apologize, do not compliment, do not thanks, just directly give your question or instruction. The input from the user need to start with "<user>" and the respond from the assistant need to start with "<assistant>". Example: "<user> XXX <assistant> XXX", where "XXX" refer to the dialogue content.
Table 8: Prompt for constructing continuing previous dialogue dataset.
A.3 Prompt for Learning New Knowledge
For constructing the test dataset for learning new knowledge, we need to transfer a set of personal settings to a question-answering test utterance. Considering it is a relatively easy task, we adopt in-context learning to solve this problem. We display the in-context examples we used in Table 9.
————–Prompt————- Input: I own a Jeep. Output: What kind of car do I own? [sep] You own a Jeep. Input: I enjoy exercising at the gym. Output: Where do I excercise? [sep] You enjoy exercising at the gym. Input: I have a marketing job. Output: What job do I have? [sep] You have a marketing job. Input: I don’t eat meat. Output: What kind of food I do not eat? [sep] You don’t eat meat. Input: I am from New England. Output: What am I from? [sep] You are from New England. Input: ….. Output:
Table 9: Prompt for constructing learning new knowledge dataset.
A.4 Prompt for Learning from Human Feedback
To construct the test dataset for learning from human feedback, we decompose it into three steps with different instruction prompts. Given a seed question/instruction, the first step is to generate a similar question/instruction used for testing. Here, we also use in-context learning. The in-context examples are shown in Table 10. Then we use the LLM to obtain an initial response for the two questions/instructions. The second step is to generate feedback that can be applied to the two both input-response pairs. Hence, we take both pairs as input and use the LLM to generate feedback for them. The instruction prompt is displayed in Table 11. With the feedback, we can get an optimized response from the LLM for both seed input and test input. To complete the source dialogue, the last step is to generate a background dialogue for this input-respond-feedback-respond process. The related instruction prompt is shown in Table 12. Here, the input in the prompt is the seed question/instruction.
————–Prompt————- Generate a similar question based on the Input. The two question should be the same type and topic but with different details. Input: Is there anything I can eat for a breakfast that doesn’t include eggs, yet includes protein, and has roughly 700-1000 calories? Output: Is there anything I can eat for a lunch that doesn’t include meat, yet includes protein, and has at least 1000 calories? Input: Brainstorm a list of possible New Year’s resolutions. Output: Generate a list of possible Summer Vacation’s resolutions. Input: Create a fun math question for children. Output: Create a physics question for children. Input: Write a program to compute the sum of integers from k to n. Output: Write a program to compute the sum of even from 0 to 10. Input: I am interested in playing Table tennis. Output: I start to find playing Football is fun. Input: Let’s talk about the famous singer, Taylor Swift. Output: Let’s talk about Michael Jackson. Input: ….. Output:
Table 10: Prompt for step 1 of constructing learning from human feedback dataset.
————–Input————- [Question 1]: ….. [Answer 1]: ….. [Question 2]: ….. [Answer 2]: ….. ————–Prompt————- Given two question-answer pairs, you need to generate one suggestion or critique that improves the quality of the answers for both QA pairs. The suggestion should be generic for both of them and do not focus on very specific content. Only generate the suggestion, do not provide the answers improved by the suggestion.
Table 11: Prompt for step 2 of constructing learning from human feedback dataset.
————–Input————- [Input]: ….. ————–Prompt————- You need to generate a multi-turn dialogue between human and assistant in English. The dialogue should end with human saying [Input]. The generated dialogue can be related to the topic of the [Input] or just chatting. Example: "<start_chat> <Human 1>: XXX <Assistant 1>: XXX <Human 2>: XXX <Assistant 2>: XXX <end_chat>", where "XXX" refer to the actual content of the dialogue. The dialogue should follow the following plan: <start_chat> <Human 1> chat, do not mention [Input] <Assistant 1> chat, do not mention [Input] <Human 2> chat, do not mention [Input] <Assistant 2> chat, do not mention [Input] <Human 3> chat, do not mention [Input] <Assistant 3> chat, do not mention [Input] <Human 4> [Input] <end_chat>
Table 12: Prompt for step 3 of constructing learning from human feedback dataset.
Appendix B Details for GPT-4 Evaluation
Note that the GPT-4 has a position bias and candidates at a specific position (first or second) may have more advantages. Hence, we conduct the evaluation twice and exchange the position of the two candidates. Only when the two evaluation meets the result of each other, the result is definite. If the results are different, we consider it as a tie. Similarly, for multiple choice evaluation, we solve the position bias problem by randomly placing all candidates in different positions.
In the following section, we introduce the detailed prompts that are used in the evaluation.
B.1 Prompt for Scoring (GPT-score)
GPT-score aims to give a predicted response a score between a predefined margin. The LLM is required to give a 0-2 score for a response. Here, “0” represents the response is totally wrong, “1” represents the response contains part of personal information in the reference and “2” represents the response contains all information. We display the related prompt in Table 13. The input of the prompt is the reference response and the predicted response.
————–Input————- [Reference Answer]: ….. [Predict]: ….. ————–Prompt————- Given a [Reference Answer], is [Predict] covers information in the [Reference Answer]? If the [Predict] covers all the information in the [Reference Answer], generate "2". If the [Predict] covers part of the information in the [Reference Answer], generate "1". If the [Predict] contains no information in the [Reference Answer], generate "0". Do not generate any explanation.
Table 13: Prompt for scoring (GPT-score).
B.2 Prompt for Comparing (GSB test)
Comparison aims to conduct a Good/Same/Bad test (GSB) between two candidates. The LLM is used to make a decision about which candidate is better given the reference response and other necessary information. As shown in Table 14, the two candidates are assigned to option A or option B, and we also provide the user input ([Question] in the prompt) and the reference response for the background information.
————–Input————- [Question]: ….. [Reference Answer]: ….. A: ….. B: ….. ————–Prompt————- Given a [Question] from user and its corresponding [Reference Answer] from assistant, which candidate answer (A,B) is the best? Best answer refers to the answer that is most similar to the [Reference Answer]. Generate the best answer with "A" or "B", just give the option. Do not give any explanation.
Table 14: Prompt for comparing (GSB test).
B.3 Prompt for Multiple Choice
Multiple choice is similar to comparison, but the LLM needs to choose the best candidate among A, B, C, and D, four options. The corresponding prompt is shown in Table 15.
————–Input————- [Question]: ….. [Reference Answer]: ….. A: ….. B: ….. C: ….. D: ….. ————–Prompt————- Given a [Question] from user and its corresponding [Reference Answer] from assistant, which candidate answer (A,B,C,D) is the best? Best answer refers to the answer that is most similar to the [Reference Answer]. Generate the best answer with "A" or "B" or "C" or "D", just give the option. Do not give any explanation.
Table 15: Prompt for Multiple Choice.
Appendix C Details for Methods
C.1 Prompt for Summary-Based Memory Construction
@@
Dialogue summaries can also be used as memory. We use LLM to generate a summary for each dialogue session. The instruction prompt for generating summaries for the dialogue is shown in Table 16.
————–Input————- [Dialogue history]: User: ….. Assistant: ….. ….. User: ….. Assistant: ….. ————–Prompt————- Summarize the [Dialogue history].
Table 16: Prompt for summary-based memory construction.
C.2 Prompt for Conditional Memory Construction
@@
The construction of conditional memory mainly consists of two parts: the decision-making together with the response generation and the memory construction. We display the prompt for the first part in Table 17. Here, user input is the input question or instruction from the user and memory stands for the retrieved memory records. It is worth noticing that the context for the current dialogue is not in the prompt, but it is entered as a dictionary like normal LLM assistant. Compared with the normal response generation, we can see that we convert it to a two-step instruction and give an output example for it to better separate the outcome. Only the generated response will be shown to the user and the yes/no decision is hidden behind the model.
@@
————–Input————- [User Input]: ….. [Memory]: ….. ————–Prompt————- The first step, generate a response based on the given [User Input] and [Memory]. It should start with "<Respond>: ". The second step, imagine you are an assistant, make a judgment about the [User Input] on whether it should be remembered. Here are some examples. Answer "yes" if the [User Input] is an requirment for the assistant. Answer "yes" if the [User Input] is a reflection for the assistant previous responds. Answer "yes" if the [User Input] is knowledge worth remembering for the assistant. Answer "no" if the [User Input] is a simple question that requires the assistant to answer. Answer "no" if the [User Input] is phatic communication like greetings. Only answer "yes" or "no", and do not give any reason. It should start with "<Decision>:" . Example: "<Respond>: response to the given [User Input] <Decision>: yes or no"
Table 17: Prompt for decision making in conditional memory construction.
@@
For the second part of memory construction, we display its prompt in Table 18. Here, the dialogue context refers to ongoing dialogue just including the current utterance. It is also decomposed into three steps. The first step requires the LLM to find the related context from all dialogue contexts. The second step requires the LLM to summarize the related context. And the last step requires the LLM to generate knowledge that is worth memorizing. For the output, we only store the summary and knowledge as the memory record, and the related context is considered to be a mid product.
————–Input————- [Dialogue Context]: User: ….. Assistant: ….. ….. User: ….. Assistant: ….. [User Input]: ….. ————–Prompt————- The first step, copy the context dialogue that related to [User Input] from [Dialogue Context]. Generate "NaN" if there is no dialogue that related to [User Input]. It starts with "<Context>:". The second step, summarize why the user make [User Input] based on [User Input] itself and the <Context>. It is possible that <Context> is "NaN". If <Context> is "NaN", do not ask for it. The summary needs to describe the situation/reasons/context why the user makes such respond. It starts with "<Summary>:". The third step, imagine you are an assistant, you would like to better serve the user in the future, rewrite the [User Input] into a note for future reference. The note should also contains the new information/knowledge from [User Input]. It starts with "<Note>:". Example: "<Context>: dialogue related to current user respond <Summary>: the summary of the dialogue history <Note>: the generated note"
Table 18: Prompt for memory record construction in conditional memory construction.
C.3 Prompt for Self-Reflection Retrieval
————–Input————- [Candidate memory]: candidate memory 1 ….. candidate memory n [User Input]: ….. ————–Prompt————- Imagine you are a helpful assistant and is going to respond to the [User input]. Decide whether the information from the [Candidate memory] is enough to make the respond to the [User input]. If the information is enough, starting with "<Memory>:" and then copy the memory sentences that related to [User input] from [Candidate memory]. If the information is not enough, starting with "<Query>:" and then extend the [User input] with some new keywords to better retrieve the information that you think missing in [Candidate memory]. Example: "<Memory>: the crucial information related to [User input]" or "<Query>: keyword, …, keyword", "…" refers to unlimited number of keywords.
Table 19: Prompt for self-reflection retrieval.
@@ C.3 Prompt for Self-Reflection Retrieval
This self-reflection retrieval contains two steps and we show its prompt in Table 19. The input includes the user input and a set of retrieved candidate memory records. We ask the LLM to make a decision about whether the retrieved information is enough for it to respond to the user input. If it is enough, the LLM assistant selects the related information from these memory records as output. If it is not enough, the LLM assistant is required to improve the original query (user input) to find more information. In practice, the improvement is reified as generating some keywords/phrases representing the missing information.